Status of virgo a short update
This presentation is the property of its rightful owner.
Sponsored Links
1 / 10

Status of Virgo a short update PowerPoint PPT Presentation


  • 77 Views
  • Uploaded on
  • Presentation posted in: General

Status of Virgo a short update. Michele Punturo. A bit of History. July 29 – August 13 2005 C6 run Duty cycle: Locking 89%, Science Mode 86% September 14-19 2005 C7 run Duty cycle: Locking 70%, Science Mode 65% 19 September 2005

Download Presentation

Status of Virgo a short update

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Status of virgo a short update

Status of Virgoa short update

Michele Punturo


A bit of history

A bit of History

  • July 29 – August 13 2005

    • C6 run

      • Duty cycle: Locking 89%, Science Mode 86%

  • September 14-19 2005

    • C7 run

      • Duty cycle: Locking 70%, Science Mode 65%

  • 19 September 2005

    • Shutdown of the interferometer and start of the ISYS upgrade


Upgrade motivations 1 2

Upgrade motivations 1/2

  • Large frequency noise in the ITF

    • When the power recycling mirror is aligned large fringes are seen

      • Cavity effect between the IMC and the PR

      • Due to the large amount of light scattered in the IMC

      • The insertion of a Faraday Isolator at the IMC output was needed to interrupt the light path from the ITF to the IMC

      • Locking was impossible in this condition

        • Temporary solution adopted decreasing by a factor 10 the light injected in the IT

          • Increase of shot noise

          • Increase of electronic noise


Upgrade motivations 2 2

Upgrade motivations 2/2

  • Structures at ~40, ~54, ~80 and 150 Hz

  • More evident when the sensitivity improves

  • Many of them were due to the modes of the IB

    • Visible on the spectrum because of a mismatch between fmod and IMC length

    • Excited by the control noise

C6

C7

(beginning)


Upgrades

LB

B8

WB

Upgrades

Q8 1

IMC

Laser

Q8 2

WE

B2

L=3km

Q2 1

NB

WI

PR

BS

L=5.6m

L=3km

B7

Q2 2

L=6m

L=6.4m

B2

NI

NE

IMC_D1T

Q8 2

RFC_DT

OB

Q8 1

IB

RFC

OMC

Q1 2

B1p

EB

B1s

Q1 1

B1p

B5


New ib

New IB


New pr

New PR

PR transfer function

  • Internal modes of the composite PR were affecting the locking of the ITF

  • Standard Virgo size PR mirror installed


Others upgrade

Others upgrade

  • Other changes:

    • New Brewster window (IT/PR link)

    • New auto alignment system for the Input Mode Cleaner

    • Laser lab work

    • Update the layout of the benches

    • New injection tower flange

    • Cabling and electronic rack displacement

    • Laser electronic upgrade

    • Environmental Monitoring

    • Linear alignment hardware upgrade

    • New and low noise RF generator

    • DAQ upgrade


Current status

Current status

  • The suspension in tower of the main components has been completed the 18th of November

  • The ISYS towers have been closed the 19/12/2005 after a long integration activity (alignment, local control reactivation,…)

    • Vacuum

  • Failure in a picomotor: tower opened again the 3/01/2006

  • Locking of the IMC on 00 mode the 05/01/2006

  • 11/01/2006:

    • Starting of the IMC autoalignment

      • Still failures in the picomotors

  • Light start to be injected in the ITF


Noise budget in c7

Noise budget in C7


  • Login